GIVEN THE FUNCTIONS F(x)=5X+7 FIND EACH OF THE FOLLOWING
F(4), F(-4), F(0)

F(4)=___________
F(-4)=__________
F(0)=______________
SIMPLFY YOUR ANSWER . TYPE AND INTEGER OR A FRACTION

Answer by Theo(13342)

function is f(x) = 5x + 7

to find f(4), substitute 4 for x to get:

f(4) = 5*4 + 7 = 20 + 7 = 27

to find f(-4), substitute -4 for x to get:

f(-4) = 5*(-4) + 7 = -20 + 7 = -13

to find f(0), substitute 0 for x to get:

f(0) = 5*0 + 7 = 7